TODOs

  • bgvdata package
  • Try as hard to finish CryoBLMs analysis during August
  • WinCC - SiPMs current & temperatures
  • Picoscope scripts for trigger monitoring - Done
    • Write more info for this
  • Contacts
    • [Didier] : Gas pressure operation issues & information
    • [Colin] : Resetting of tunnels after restart of bgvgw machine
    • [Colin] : Resetting of time sync after restart of bgvopc machine

[25 / 07 / 2016]

  • BGV LHC Intervension
    • Problems with multiple now LV channels which are tripping.
    • Swapping LV CAEN cards did not result in anything good.
    • Cooling is almost done, seems like working nicely, tests OK
      • To access / control the chiller, log to bgvopc / Documents a shortcut for remote access / UI is intuitive.
      • Temperature down to 12 C for now. Insulation will be needed for more cooling.

  • Progressed a bit with the "digital scope" project - need to define a standard way of doing things.

[29 / 07 / 2016]

  • MD Week
    • Today some separate beams - check if needed to start DAQ for checking trigger with each beam separately.
    • No input from Quentin, no DAQ
  • Finishing the architecture for the data analysis / display.
    • Play-Stop-Delay , Filters, and a lot of changes in the PyQtGraph plotting module.
Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2016-08-01 - AndreasAlexopoulos
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback